# Lifting the Veil on the Large Language Model Supply Chain: Composition, Risks, and Mitigations

### v2

@@ 3. Methodology
In this section, we elaborate the methodology for our study. The overview of our methodology is presented in Figure 3.
@@ 3.1. Literature Collection
We focused our literature collection on academic literature and online resources. It helps to cover a comprehensive scope for our study. On the one hand, academic literature provides a rigorous understanding of the risks and mitigations in the LLM supply chain. On the other hand, online resources provide real-time, validated observations and practical examples that reflect current real-world practices.
@@ 3.1.1. Academic Literature Collection
The academic literature collection consists of two phases. First, we collected the initial papers set by querying with keywords that are central to our topic. Second, we expanded the paper set using the snowballing technique, incorporating both the references cited by the initial set and the papers that cite them.
Initial Paper Set. Two PhD students with over five years of experience in software supply chain management and two PhD students with over three years in LLM security have engaged in the process. Specifically, they first collected keywords from three domains. i.e., supply chain security (e.g., security, attack, malware, threat, vulnerabilities), LLM security (e.g., data poisoning, data provenance, model poisoning) and LLM DevOps (e.g., data validation, model validation, model evaluation). The remaining participants cross-checked the keywords, and together they compiled the final comprehensive set. Afterwards, we constructed queries based on the following rules using the advanced search interface on Google Scholar.
@@
Figure 3. Approach Overview of our Literature Review
Furthermore, we excluded papers that do not align with the goal of our study. First, we filter out papers published before 2018. Second, we exclude short papers with fewer than 2 pages. Finally, we manually review the remaining papers to assess their relevance to the large language model supply chain. Consequently, this process leaves us with 127 papers.
Snowballed Paper Set. We identified cited and citing papers from the initial set, with each new group of relevant papers forming the basis for the next round of reviews. This iterative process continued until we reached a point where no new relevant papers were found. Ultimately, this process led us to a final set of 313 papers.
@@ 3.1.2. Online Resource Collection
In addition to collecting scientific literature, we have also gathered online resources which often include real-world attacks or the latest knowledge not yet covered by scientific literature.
Table 1 lists the websites where we collected online resources. We chose these websites for their wide coverage, reputation in the field of cybersecurity, professionalism, timely updates, and authoritative information. On the one hand, we leverage the same keywords used in the academic literature collection on the websites containing querying interfaces. e.g., Google. On the other hand, we manually inspect the websites to collect the relevant news or events.
Table 1. Websites of Online Resource
Websites Link
Google https://www.google.com/
MITRE ALTAS https://atlas.mitre.org/
GitHub Blog https://github.blog/
JFrog https://jfrog.com/blog
Sonatype https://www.sonatype.com/blog
Synk https://snyk.io/blog/
Tencent Security https://security.tencent.com/index.php/blog
Hacker News https://thehackernews.com/
Checkmarx https://checkmarx.com/blog/
Check Point https://blog.checkpoint.com/
BleepingComputer https://www.bleepingcomputer.com/news/security
The Register https://www.theregister.com/security
Security Intelligence https://securityintelligence.com/news/
heise https://securityintelligence.com/news/
SC Magazine https://www.scmagazine.com/
3.2. Literature Analysis and Taxonomy Construction
We manually inspected and categorized various elements of large language model risks and mitigations from the literature. First, we used open card sorting in our analysis. We began by thoroughly reviewing the literature set, and identified and highlighted terms, concepts, techniques, and strategies related to the large language model risks and mitigations.
We performed hierarchical tree structure construction, where we connected the identified categories into a structured, hierarchical framework that reflects the relationships between different risks and mitigation elements. We organize these categories into tree structures where the root represents the risks, stakeholders, and mitigations. To ensure accuracy and comprehensiveness, we refine our summarized structures according to the following criteria with multiple iterations.
@@ 3.2. Literature Analysis and Taxonomy Construction
  • Relevance. Ensure that each node reflects the latest risks and mitigations by reviewing recent research and industry standards (e.g., CVE, CWE, MITRE ATT&CK).
@@
Each node in the structures is assessed by the four professionals independently. The structures are refined and re-evaluated iteratively until all professionals reach a final agreement.
4. Risks of Large Language Model Supply Chain
In this section, we summarize the risks in the LLM supply chain through illustrative risk scenarios. The scenarios are presented as risky stakeholders performing risky actions on specific supply chain components, which in turn lead to distinct risk types. We first discuss the risky stakeholders. Then, we study the risk types. Next, we propose the taxonomy of risky actions in the large language model supply chain. The risky stakeholders are detailed in the Section 2.2.1 and the supply chain components are covered in the Section 2.2. The risky actions are illustrated in Section 4.2 and the risk types are illustrated in Section 4.3. We present the risk scenarios regarding Risky Contributors, Risky Consumers, Risky Administrators and Risky Users in Figure 5, 6, 8 and 7, respectively.
4.1. Risky Stakeholders
We discuss the risky stakeholders in terms of risky contributors, risky consumers, risky administrators and risky users.
  • Risky Contributors.
    Camouflage Contributors. These contributors attempt to blend in with legitimate contributors, possibly hiding their true intentions.
    Attack Contributors. This group actively seeks to compromise the system by injecting malicious code, data, or models into the supply chain.
    Being Benign Contributors. Contributors in this category introduce vulnerabilities unintentionally, either through poor coding practices, lack of maintenance, or other factors that increase system risk.
  • Risky Consumers.
    Being Middleman. Entities that act as intermediaries, facilitating the exchange or integration of supply chain components between contributors and other consumers.
    Being Artifact Consumer. These stakeholders use artifact (e.g., data) for training, analytics, or other purposes.
    Being Platform Consumer. Entities that rely on platforms built with supply chain components, such as development or deployment environments.
  • Risky Administrators.
    Attack Administrators. Attackers can hack into an administrator’s account, causing severe consequences given the elevated privileges and control administrators hold over systems.
    Being benign administrators. Administrators who despite having good intentions, inadvertently introduce risks into the system through misconfigurations or other unintentional actions.
  • Risky Users. Users can be risky if their purposes are to exploit LLM application vulnerabilities, steal models, conduct jailbreaking or inference attacks, perform excessive access, or pollute the data using user feedbacks.
Figure 4. Risk Types
@@ 4.2. Risk Types
The LLM supply chain risks are classified into four main types. i.e., Security Risks, Privacy Risks, Delivery Risks, and Legal Risks. Figure 4 provides a comprehensive overview of the LLM supply chain risks across security, privacy, operational, and legal domains.
  • Security Risks (SR) include potential threats such as hacker use (SR1), system compromise (SR2), and the emergence of abnormal content (SR3) that could disrupt systems or compromise integrity.
@@ 4.3. Taxonomy of Risky Actions
We present the taxonomy of LLM supply chain risks with regard to the six composing elements in the LLM supply chain.
@@
Figure 5. Risky Actions, Supply Chain Components and Risk Types Introduced by Risky Contributors
4.3.1. Risks on Code
The code, including packages, frameworks, and plugins, have become essential at every stage of LLM application development. However, using these libraries comes with potential risks.
@@ 4.3.1. Risks on Code
Implant Malicious Code. Attackers can develop malicious code and attempt to implant it through various methods, finally integrating it into the LLM supply chain. For instance, Zhao et al. (Zhao et al., 2024b) revealed 9 malicious dataset loading scripts on Hugging Face, and perform root cause analysis of vulnerable formats. It can cause the risk of malicious use (LR2).
@@
  • Dependency Confusion. A dependency confusion attack is a type of supply chain attack where an attacker publishes a malicious package to a public package repository with the same name as a popular private package. The package manager may unknowingly download the malicious package from the public repository instead of the intended private repository (PyTorch, 2022).
  • Typosquatting. Using typosquatting or combosquatting, attackers upload malicious packages with names that are nearly identical to popular legitimate packages. A minor typographical error during package installation can result in the inclusion of malicious code in the application. For example, PyPI has suffered from numerous malicious typosquatting packages. These malicious packages can lead to the leakage of Personally Identifiable Information (PII) and compromise the integrity and security of systems (CheckPoint, 2024). Rokon et al. (Rokon et al., 2020) identified over 7,500 instances of malware hosted on GitHub. Attackers may also employ more covert methods, such as spreading malware through forks (Cao and Dolan-Gavitt, 2022).
@@
  • Compromising Legitimate Libraries. Attackers may target legitimate libraries already integrated into the LLM supply chain, leveraging their established presence and substantial user base. Attackers can conceal their malicious code within pull requests, subtly injecting harmful elements into the codebase. In addition, attackers may take control of the code from a previous maintainer and release new malicious versions (II, 2018).
@@
  • Sharing Cloud Environments. Attackers can inject malicious code into cloud environments and make it publicly available to all cloud users. The rise in popularity of containers (e.g., Docker) has accelerated the trend of deploying applications in cloud environments. However, it in turn has become a hugely popular attack vector recently. The research team discovered that over a thousand Docker container images were found hiding malicious content (Chierici, 2022). Tomar et al. (Tomar et al., 2020) proposed a taxonomy of potential attacks at the container layer, revealing different types of threats that can arise within containerized environments.
4.3.2. Risks on Dataset
The datasets play a fundamental role, providing the extensive data required for model training, model refining and prompt generation.
@@ 4.3.2. Risks on Dataset
Implant Poisoned Data. Attackers can create poisoned data and implant it through various methods, with the aim of having it used during critical stages of building the LLM supply chain. It can cause the risk of data leakage (PR1) and malicious use (LR2).
  • Data Poisoning. Attackers can manipulate the data for malicious purposes, and then upload it to public platforms (e.g., Kaggle) (Ma et al., 2019). Given that training large language models requires vast amounts of data, it is highly likely that these poisoned data sets will be used by LLM model developers. If the poisoned data is unintentionally leveraged into the LLM supply chain, it would result in the model performance degradation, particularly in its ability to generate the outputs accurately and efficiently (Shu et al., 2023; Wan et al., 2022; Shan et al., 2024). Alexander et al. (Wan et al., 2023) demonstrated that with only 100 poisoning examples, LLM produces consistent negative results or vulnerability output across hundreds of tasks.
  • RAG Poisoning. Retrieval-augmented generation (RAG) is widely incorporated in the LLM supply chain to provide the LLM with sufficient contextual knowledge. Similarly, the knowledge database can also be poisoned, where it can be manipulated the guide the LLM to generate specific, attacker-chosen responses to particular questions. Zou et al. (Zou et al., 2024) proposed PoisonedRAG, which specifically focuses on injecting poisoned texts into knowledge databases to steer language models to respond with predefined answers to questions selected by the attacker. Wang et al. introduce the concept of Poisoned-LangChain (Wang et al., 2024a), a novel indirect attack that exploits a poisoned external knowledge base.
4.3.3. Risks on Models
The open-source models provide accessibility, flexibility, and transparency, allowing developers to modify, customize, and integrate them into various projects.
@@ 4.3.3. Risks on Models
Implant Model Backdoors. Attackers can manipulate publicly accessible models to alter their behavior for malicious purposes, ultimately affecting downstream users who rely on the compromised models. LLM may be composed of multiple components, including vocabularies, tokenizers, embedding layers, and auxiliary models (e.g., CLIP-ViT), etc (Liu et al., 2024; Radford et al., 2019; Chen et al., 2023). It can cause hacker use (SR1), system compromise (SR2), emergence of abnormal content (SR3), privacy risk (PR), data leakage (PR1), personally identifiable information (PII) leakage (PR2), model stealing (PR3), code leakage (PR4), and prompt leakage (PR5).
  • Component Dependency Attack. Attackers can manipulate the contents or structures of the composing models of LLMs, thereby conducting component dependency attacks. Huang et al.(Huang et al., 2023) firstly explore how manipulating the embedding dictionary using carefully designed rules could cause the model to produce specific outputs. Additionally, Cui et al.(Cui et al., 2024a) and Shayegani et al. (Shayegani et al., 2023) demonstrated that LLMs are not robust against attacks targeting the visual models used exclusively in multi-modal LLMs.
@@
  • Backdoor attack. Backdoor attack embeds hidden backdoors in the model during the training process, allowing the compromised model to perform normally on benign samples but altered on the hidden backdoor input (Li et al., 2021; Chen et al., 2021; Pan et al., 2022). Yao et al. and Cai et al. (Cai et al., 2022; Yao et al., 2024b) studied the vulnerability of prompt learning algorithms commonly used in LLM to backdoor attacks.
@@
Inference Attack. Inference attacks are types of privacy attack targeting data, where the responses of LLM models may inadvertently leak sensitive information, leading to unauthorized access, intellectual property theft, and privacy breaches. They include attribute inference attacks and membership inference attacks. Attribute inference attacks involve deducing sensitive information (e.g., race, gender, and sexual orientation) from the behavior or responses of LLM models, even if such information is not explicitly included in the training data. Robin et al. (Staab et al., 2023) and Pan et al. (Pan et al., 2020) conducted systematic studies on various advanced large language models, which have shown that LLMs can accurately infer sensitive data such as identity, genome, healthcare, and location information without any prior knowledge. Membership inference attacks aim to predict whether a data sample was included in the training data of LLM, undermining the trust between data providers and users. Mattern et al. (Mattern et al., 2023) explored the significant threat of membership inference in the real-world using neighborhood comparison. It can cause data leakage (PR1), personally identifiable information (PII) leakage (PR2), model stealing (PR3), code leakage (PR4), and prompt leakage (PR5).

### v3

@@ 3. Methodology
Following the systematic literature review guidelines (Garousi and Mäntylä, 2016), we adopt a scientific and systematic approach for literature collection (see Sec. 3.1), then we performed a taxonomy analysis (see Sec. 3.2). An overview of our review process is presented in Figure 2.
Figure 2. Approach Overview of our Literature Review
@@ 3.1. Literature Collection
We review the academic literature and online resources to collect an extensive list of risks and mitigations about LLM supply chain.
@@ 3.1.1. Academic Literature Collection
We define the temporal scope of our survey to span from the emergence of the earliest relevant publications in this field to the most recent contributions, specifically covering the period from January 2018 to December 2024 (i.e., around 7 years). Two PhD students with over five years of experience in software supply chain management and two PhD students with over three years in LLM security have engaged in this process. The detailed steps are as follows.
Initial Paper Querying. We identify potentially relevant papers by conducting searches across major electronic scholarly databases. Specifically, we select ACM Digital Library [^1] https://dl.acm.org/, IEEE Xplore [^2] https://ieeexplore.ieee.org/Xplore/home.jsp, DBLP [^3] https://dblp.org/ and Semantic Scholar [^4] https://www.semanticscholar.org/me/research as our primary data sources, which are popular bibliography databases containing a comprehensive list of research venues in computer science. Then, we define the keywords presented in Listing 1 from three domains, i.e., supply chain security (e.g., security, attack, malware, threat, vulnerabilities), LLM security (e.g., data poisoning, data provenance, model poisoning) and LLM DevOps (e.g., data validation, model validation, model evaluation).
Listing 1: Search Keywords for Literature Collection
@@
Furthermore, we perform a manual assessment to determine the relevance of each candidate paper to the topic of LLM supply chain, excluding those that do not align with the goal of our study. Besides, we exclude short papers with fewer than 2 pages, as well as survey paper papers or summary papers, which are separately discussed in Sec. 2 for a comparison with our work. Consequently, we obtain an initial set of 127 papers.
Paper Snowballing. To mitigate the risk of omitting relevant papers, we apply both backward and forward snowballing techniques (Wohlin, 2014) on the initial paper set. In the backward snowballing phase, we check the reference lists of these papers to discover new candidates. In forward snowballing phase, we use Google Scholar [^5] https://scholar.google.com/ to identify papers that cite the initial set. This iterative process is repeated until no further relevant papers are found. All newly identified candidate papers are also manually assessed for relevance, resulting in a final set of 238 papers.
@@ 3.1.2. Online Resource Collection
In addition to collecting papers in the academic literature, we also include online resources, e.g., blog posts and incident reports, which often include real-world attacks or the latest knowledge not yet covered by academic literature. These websites are selected based on their broad coverage, strong reputation in the cybersecurity community, professional content, timely updates, and authoritative reporting. Following a procedure similar to that used for academic literature collection, on the one hand, we leverage the same keywords used in the academic literature collection on the websites containing querying interfaces. On the other hand, we manually inspect the websites to collect the relevant news or events, resulting in 31 reports related to our topic.
3.2. Taxonomy Construction
We download all the relevant resources and conduct a full-text analysis to inspect and categorize various elements of large language model risks and mitigations from the literature. First of all, we use open card sorting in our analysis to highlight terms, concepts, techniques, and approaches related to the large language model risks and mitigations. Then, we perform the hierarchical tree structure construction, where we connect the identified elements into a structured, hierarchical framework that reflects the relationships between different aspects of the LLM supply chain. We organize tree structures where the root represents the risks, stakeholders, and mitigations. At last, we refine our summarized structures according to the following criteria with multiple iterations.
  • Relevance. Ensure that each node reflects the latest risks and mitigations by reviewing recent research and industry standards.
@@ 3.2. Taxonomy Construction
Each node in the structures is assessed by the four professionals independently. The structures are refined and re-evaluated iteratively until all professionals reach a final agreement. Furthermore, the label is left blank if the literature does not cover specific aspects (e.g., mitigation).
4. Composition of Large Language Model Supply Chain (RQ1)
We present an overview of the large language model supply chain in Figure 1. In general, we identify four types of stakeholders, interacting with specific components of the large language model supply chain.
4.1. Stakeholders
We summarize four roles of stakeholders in the LLM supply chain. i.e., contributors, consumers, administrators, and user.
Contributors. Person who engaged in development activities and sharing artifacts, such as data, models, prompts, or third-party libraries, etc. Platform Contributors are responsible for contributing to hosting platforms, hardware platforms, and serving platforms (i.e., Amazon Web Services (Services, 2024)) and providing stable, isolated, and managed environments. Artifact Contributors process, and upload various forms of data, code, prompts, and models to the hosting platforms. They can be data engineers, model developers, prompt engineers, and library developers. Toolchain Contributors are responsible for developing, maintaining, and improving the tools used throughout the model lifecycle. e.g., Jupyter, TensorFlow for model development, Jenkins, GitHub Actions in CI/CD.
Consumers. We distinguish consumers depending on their consuming components, including platform consumers, artifact consumers, toolchain consumers, and app consumers. Platform Consumers are entities that use the platform infrastructure for various purposes, such as accessing hosting services, utilizing computational resources, or interacting with serving platforms for deploying and running models and applications. Artifact Consumers are those who utilize the artifacts produced and shared within the platform. This includes data sets, code, prompts, and models. Toolchain Consumers are individuals or teams that leverage the toolchain components for developing, deploying, and maintaining models. They use model development, CI/CD, and deployment tools to streamline the lifecycle of models and applications. App Consumers are end-users or other systems that interact with applications built using the platform’s resources and artifacts.
Figure 3. Distribution of Components that Covered in our Collected Literature
Administrators. We summarize administrators as platform administrators and apps administrators. Platform Administrators represent a group of people who manage and oversee the infrastructure and another group who manage repositories of artifacts. They hold the privileges to merge pull requests, manage versioning, and onboard new contributors. Apps Administrator is responsible for managing the lifecycle of applications that rely on the artifacts and models within the platform. This includes deploying, configuring, and monitoring applications.
4.2. Components Analysis
We summarize four components, including artifact, toolchain, platform, and LLM applications (apps). Figure 3 illustrates the distribution of supply chain components covered in the surveyed literature. Notably, the majority of the works (218 papers, 71.9%) focus on risks associated with artifacts, followed by applications (66 papers, 21.8%). In contrast, significantly fewer studies address risks related to the platform (11 papers, 3.6%) and toolchain (8 papers, 2.6%). Given that the platform and toolchain serve as foundational and widely adopted components in the LLM supply chain, their limited coverage in existing research underscores a critical gap and highlights the need for greater research attention.
4.2.1. Artifacts
The LLM supply chain artifacts include data, model, prompts and code.
Data. The data can be divided as the corpus for pre-training or fine-tuning, and the knowledge database for retrieval augmented generation.
Model. The model represents the large language model that is trained and fine-tuned. The model may be adjusted or modified by splitting, merging, serializing, or deserializing components as needed to optimize performance.
Prompt. The prompt is crafted to instruct the model effectively, guiding it on what information to retrieve, process, or generate.
Code. The code encompasses the software and scripts utilized for data processing, model training, and prompt refinement. This includes code that directly invokes relevant library APIs, as well as the libraries themselves as dependencies.
4.2.2. Platforms
Hosting Platform. These platforms facilitate the storage, and sharing of software packages, models, data, and related resources required for LLM development and deployment. e.g., Data Hosting Platforms (e.g., Kaggle (LLC, 2024)), Model Hosting Platforms (e.g., Hugging Face (Face, 2024)), Repository Hosting Platforms (e.g., GitHub (GitHub, 2024a)), Package Registries (e.g., PyPI (Foundation, 2024)), Plugin Marketplaces (e.g., Visual Studio Marketplace (Microsoft, 2024)).
Hardware Platform. These platforms consist of specialized hardware designed to optimize the performance of LLMs.
Serving Platform. These platforms offer cloud-based services to support the training, deployment, and maintenance of LLMs.
4.2.3. Toolchain
The development and deployment of large language models (LLMs) involve two connected toolchains: model development toolchain and model deployment toolchain.
Model Development Toolchain. Development toolchain encompasses the tools and workflows used to train, refine, and optimize the model. Key components include training frameworks, model serialization and versioning, model conversion and model quantization tools.
Model Deployment Toolchain. Deployment toolchain focuses on integrating the trained model into production environments throughout the model’s lifecycle in production. It automates deployment, monitoring, and real-time updates. Core components include CI/CD systems, which ensure seamless integration and continuous updates in live environments, and cloud services or containerization tools (e.g., Docker) that provide the deployment environment.
4.2.4. Applications
Summary. We examined four key components in the LLM supply chain: artifacts, platforms, toolchains, and applications. These components can be further broken down into four types of artifacts, three categories of platforms, three toolchain types, and four applications. While most studies focus on risks related to artifacts and applications, the limited attention to the widely adopted and foundational components, i.e., platform and toolchain, reveals a critical research gap in the LLM supply chain.
4.3. Composition Relation Discussion
We discuss the composition of relations within the LLM supply chain from the perspectives of explicit and implicit relations. The relation is derived from components across artifacts, toolchains, applications, platforms, and other related elements.
4.3.1. Explicit Relations
Explicit relations are those that can be directly identified from artifacts such as source code, configuration files (e.g., package manager configuration files), and documentation.
Invocation Relations from Source Code. Explicit dependencies are often revealed through import statements or direct method calls. These include class or method imports, method invocations, the use of imported constants, and API invocations.
Dependency Declaration in Config Files. Dependencies explicitly declared in configuration files, such as package.json in JavaScript or pom.xml in Java, provide clear indications of relationships between components. These files serve as a formal record of required libraries or modules.
Dependency Statements in Documentation. Explicit relations can also be derived from documentation, such as API references, README files, or inline comments, which explicitly describe dependencies, interactions, or usage guidelines.
4.3.2. Implicit Relations
Implicit relations are those embedded during the development and deployment of LLMs. Unlike explicit relations, they cannot be directly inferred from existing artifacts. Nevertheless, they are an inevitable and integral component of the LLM supply chain.
Training Relations from Training and Fine-tuning Dataset. Data stands as one of the most critical components in training LLMs. The direct dependencies on data are consolidated into the trained weights, which are inherently embedded within the models. While explicit patterns are directly extracted, implicit relations, though not directly extractable, can still be inferred through techniques such as membership inference attacks (Feng et al., 2024; Duan et al., 2024).
Prompt Augmentation Relations from RAG Dataset and Third-party Prompts. The RAG dataset serves as an external knowledge source, enabling LLMs to acquire a more comprehensive understanding of the queried questions. Meanwhile, third-party prompts offer detailed and tailored instructions, making it easier for LLMs to interpret and respond effectively. These elements act as supplementary knowledge sources, providing richer contextual information for LLMs. As a result, we define these relationships as prompt augmentation relations.
Artifact Reusing Relations. In traditional software development, code reuse commonly occurs through a well-known practice called code cloning. As we transition to the LLM supply chain and encounter new types of artifacts, novel forms of reuse have emerged. For example, a most notably new type is model reuse. To fulfill specific requirements, downstream users may split, merge, or convert models, reflecting more dynamic and complex reuse patterns. Moreover, datasets often undergo preprocessing as a standard part of the pipeline.
Summary. We explore both explicit and implicit relations within the LLM supply chain. Implicit relations, in particular, deserve significant attention due to their inherent complexity in inference and their pervasive presence across the LLM supply chain. These represent compositional relationships that may introduce new risks and challenges.
5. Risks of Large Language Model Supply Chain (RQ2)
We introduce the notion of risky scenarios. Then, we illustrate risk types, risky stakeholders, risky actions, respectively.
5.1. Risky Scenarios
The risky scenarios can be illustrated that a group of risky stakeholders perform some risky actions on specific supply chain components, causing certain risk types. The risky stakeholders, risky actions, supply chain components, and risk types are presented from the first column to the fourth column in Figure 6, respectively.
@@ 5.2. Risk Types
Figure 4 illustrates the distribution of risk types addressed in the surveyed literature. A majority of the studies (205 papers, 54.2%) concentrate on security risks, followed by privacy risks (91 papers, 24.1%). In contrast, legal risks (49 papers, 13.0%) and delivery risks (33 papers, 8.7%) have received comparatively less attention. This uneven distribution calls for greater attention to legal and delivery risks, which remain underexplored despite their potential impact on the LLM supply chain.
Figure 4. Distribution of Risk Types that Covered in our Collected Literature
Figure 5. Distribution of Stakeholders that Covered in our Collected Literature
  • Security Risks (SR) include potential threats such as hacker use (SR1), system compromise (SR2), and the emergence of abnormal content (SR3), i.e., manipulated content, frad, misinformation, etc that could disrupt systems or compromise integrity.
@@
5.3. Risky Stakeholders
Figure 5 illustrates the distribution of risky stakeholders addressed that are discussed in the surveyed literature. A majority of the studies (134 papers, 39.2%) concentrate on risky consumers, followed by risky users (113 papers, 33.0%), and risky contributors (91 papers, 26.6%). Few papers discuss the risk introduced by the risky administrators (4 papers, 1.2%). The results present a potential research gap in understanding and mitigating risks introduced by those who develop, deploy, or manage LLM applications.
  • Risky Contributors: (a) Malicious contributors can either camouflage legitimate contributors to gain system access or penetrate into systems by leveraging system vulnerabilities. (b) Benign contributors may unintentionally introduce vulnerabilities, either through poor coding practices, lack of maintenance, etc. (c) Volatile contributors may suspend their support under certain circumstances.
  • Risky Consumers: Attackers can play the role of artifact consumers who can (a) act as middlemen that can inject malicious content during the exchange, or (b) re-develop existing artifact for malicious use. Besides, attack can also (c) be platform consumers who can exploit platform vulnerabilities.
  • Risky Administrators: (a) Attackers can hack into an administrator’s account, causing severe consequences given the elevated privileges and control. (b) Benign administrators can inadvertently introduce risks into the system through misconfigurations or other unintentional actions.
  • Risky Users: Users whose purposes are to exploit LLM application vulnerabilities, steal models, conduct jailbreaking or inference attacks, perform excessive access, or pollute the data using user feedbacks.
@@ 5.4. Taxonomy of Risky Actions
Figure 7 illustrates the distribution of risky actions addressed in the surveyed literature. The five most frequently discussed actions are exploiting LLM application vulnerabilities (90 papers, 21.1%), implanting poisoned prompts (74 papers, 17.4%), implanting poisoned data (59 papers, 13.8%), performing model inference attacks (51 papers, 12.0%), and implanting model backdoors (33 papers, 7.7%). These actions are largely inherent to the unique characteristics of LLMs and represent threats that are distinct from those observed in traditional software supply chains. In contrast, only a few papers address actions such as degrading sustainability (4 papers, 0.9%), data leakage (4 papers, 0.9%), and discontinuation of maintenance (1 paper, 0.2%). Given the continuous operation and evolution of LLM applications, this limited attention to long-term risks may pose significant challenges in the long run.
@@
Figure 6. Risky Actions, Supply Chain Components and Risk Types Introduced by Risky Stakeholders
Figure 7. Distribution of Risky Actions that Covered in our Collected Literature
5.4.1. Risks on Artifact
The artifact includes code, data, model and prompt.
@@ 5.4.1. Risks on Artifact
Implant Malicious Code. Attackers can develop malicious code and attempt to implant it through various methods. For instance, Zhao et al. (Zhao et al., 2024b) revealed 9 malicious dataset loading scripts on Hugging Face. It can cause the risk of malicious use (LR2).
@@
  • Dependency Confusion. Attackers publish a malicious package to a public package repository with the same name as a popular private package. The package manager may unknowingly download the malicious package from the public repository instead of the intended private repository (PyTorch, 2022).
  • Typosquatting. Attackers upload malicious packages with names containing minor typos to imitate popular legitimate packages. Accidentally installing those packages can lead to security issues and compromise the integrity and security of systems (CheckPoint, 2024).
@@
  • Compromising Legitimate Libraries. Attackers may target legitimate libraries already integrated into the LLM supply chain. They can conceal their malicious code within pull requests. In addition, attackers may take control of the code from a previous maintainer and release new malicious versions (II, 2018).
@@
  • Sharing Cloud Environments. Attackers can inject malicious code into cloud environments and make it publicly available to all cloud users. It was discovered that over a thousand Docker container images were found hiding malicious content (Chierici, 2022). Tomar et al. (Tomar et al., 2020) proposed a taxonomy of potential attacks at the container layer.
@@
Implant Poisoned Data. Attackers can create poisoned data and implant it during critical stages of building the LLM supply chain. It can cause the risk of data leakage (PR1) and malicious use (LR2).
  • Training Data Poisoning. Attackers can manipulate the data for malicious purposes, and then upload it to public platforms (e.g., Kaggle) (Ma et al., 2019). Once integrated into the LLM supply chain, it would result in the model performance degradation (Carlini et al., 2024a; Shu et al., 2023; Wan et al., 2022; Shan et al., 2024; Wan et al., 2023).
  • Retrieval Data Poisoning. Retrieval-augmented generation (RAG) is widely incorporated to provide the LLM with sufficient contextual knowledge. The database can also be poisoned, guiding the LLM to generate specific, attacker-chosen responses (Zou et al., 2024; Wang et al., 2024a).
@@
Implant Model Backdoors. Attackers can manipulate publicly accessible models to alter their behavior for malicious purposes, ultimately affecting downstream users who rely on the compromised models. Generally, the models refer to various components, including the foundation model, vocabularies, tokenizers, embedding layers, and auxiliary models, etc (Liu et al., 2024; Radford et al., 2019; Chen et al., 2023). It can cause hacker use (SR1), system compromise (SR2), emergence of abnormal content (SR3), privacy risk (PR), data leakage (PR1), personally identifiable information (PII) leakage (PR2), model stealing (PR3), code leakage (PR4), and prompt leakage (PR5).
  • Component Dependency Attack. Attackers can manipulate the contents or structures of the composing models of LLMs, affecting the robustness (Shayegani et al., 2023; Cui et al., 2024a) and causing unexpected consequences(Huang et al., 2023).
@@
  • Backdoor attack. Backdoor attack embeds hidden backdoors in the model during the training process, allowing the compromised model to perform normally on benign samples but altered on the hidden backdoor input (Li et al., 2021; Chen et al., 2021; Pan et al., 2022).
@@
Perform Model Inference Attacks. Inference attacks aim to exploit LLM models, causing them to inadvertently disclose sensitive information. They include attribute inference attacks and membership inference attacks. Attribute inference attacks involve deducing sensitive information (e.g., race, gender, and sexual orientation) from the behavior or responses of LLM models, even if such information is not explicitly included in the training data . (Staab et al., 2023; Pan et al., 2020). Membership inference attacks aim to predict whether a data sample was included in the training data of LLM (Mattern et al., 2023). It can cause data leakage (PR1), personally identifiable information (PII) leakage (PR2), model stealing (PR3), code leakage (PR4), and prompt leakage (PR5).
Implant Poisoned Prompts. Attackers can conceal harmful or misleading prompts into publicly accessible repositories or datasets, which are widely shared and reused for downstream users. It can cause the performance degradation (DR1) of delivery risks.
  • Prompt injection. Prompt injections disguise malicious instructions as benign, exploiting the incapability of LLMs on distinguishing “good” and “bad” prompts (Greshake et al., 2023; Chen et al., 2024; Shayegani et al., 2023).
  • Jailbreaking. Jailbreaking attempts to bypass safety filters built in the LLMs (Willison, 2024), making the LLMs responsive to restricted or insecure content (Wei et al., 2024; Deng et al., 2024; Zou et al., 2023; Shen et al., 2023).